0.10 0sx<1 0.15 1 χ<2 0.40 2 <3 0.70 3 x <4 1 x24 7. (10 points) Let X be a discrete random variable with CDF: Fx(x)- Find P(0< X<4)
Answer #1

To calculate P(a<X<b)
where a<b. Recall, that if F is the CDF of X then F(x)=P(X≤x) for all values x. Now use this in given problem

P(a<X<b)=P(X<b)−P(X≤a)

P(0<X<4) = F(4) - F(0) = 0.70 - 0 = 0.70